John, in going to his Physics class, walks 30m, 450 N of E, then 40 m, 600 W of S and finally 50 m, straight south. Use the polygon method to determine the resultant vector 𝑅⃑ and check your result using the component method.


1
Expert's answer
2022-03-11T08:28:27-0500



Let us determine the distance OB:

"OB^2 = OA^2+AB^2 - 2OA\cdot AB\cos(60^\circ-45^\circ),\\
OB^2 = 30^2+40^2 - 2\cdot30\cdot40\cos15^\circ,\\
OB = 13.5\,\mathrm{m}."


"\dfrac{\sin\angle AOB}{AB} = \dfrac{\sin OAB}{OB}, \\
\angle AOB = 130^\circ. \\
\angle BOW = 180^\circ - 130^\circ - 45^\circ = 5^\circ."

"\angle OBW = 90^\circ -\angle BOW = 85^\circ."


"|\vec{R}|=OC -?\\
OC^2= BC^2 + OB^2 - 2BC\cdot OB\cos\angle OBW,\\
OC = 50.6\,\mathrm{m}."


"\dfrac{\sin\angle BOC}{BC} = \dfrac{\sin\angle OBW}{OC}, \;
\angle BOC = 80^\circ.\\
\angle COS = 90^\circ - (\angle BOC - \angle BOW) = 15^\circ."



Let us determine the parameters of R using the coordinate method.

"\vec{R}_x = \vec{OA}_x + \vec{AB}_x + \vec{BC}_x, \\
\vec{R}_x =30\cos45^\circ + 40\cos(180^\circ + 30^\circ) + 50\cos270^\circ, \\
\vec{R}_x = -13.4.
\\
\vec{R}_y = \vec{OA}_y + \vec{AB}_y + \vec{BC}_y, \\
\vec{R}_y =30\sin45^\circ + 40\sin(180^\circ + 30^\circ) + 50\sin270^\circ, \\
\vec{R}_y = -48.8."